# Goal Reaching with Eikonal-Constrained Hierarchical Quasimetric Reinforcement Learning

### v1

@@ 1 Introduction
In recent years, Reinforcement Learning (RL) has emerged as a powerful paradigm for solving complex decision-making tasks from experience, achieving success in domains such as game playing (mnih2015human; silver2017mastering), robotics (tang2025deep), autonomous driving (kiran2021deep; sabouni2024reinforcement; ahmad2025hierarchical), and industrial control (mirhoseini2020chip). Despite these achievements, the aforementioned solutions fundamentally rely on human-designed rewards, a process that is notoriously difficult and time consuming. Reward engineering introduces hand-crafted structures, additional hyperparameters, and assumptions about task solutions, which ultimately undermines RL’s promise to transcend human design and ability.
@@
Goal-Conditioned RL (GCRL) provides an elegant way to address this challenge by framing problems in terms of reaching arbitrary goals (schaul2015universal; andrychowicz2017hindsight). By conditioning policies on goals, GCRL offers a scalable route toward flexible, goal-directed behavior without the need for manually designed reward functions. Beyond this practical advantage, recent work has uncovered a fundamental geometric property of GCRL: the optimal goal-conditioned value function $V^{*}(s,g)$ corresponds to the length of the shortest feasible path from a state $s$ to a goal $g$, and therefore naturally defines a quasimetric (sontag1995abstract; liu2023metric).
@@
This perspective underpins Quasimetric RL (QRL) (wang2023optimal), a recently proposed GCRL algorithm that restricts value functions to quasimetric mappings, reducing the hypothesis space from arbitrary functions to a structured subset aligned with goal-reaching tasks. In this framework, GCRL is recast as an optimization problem over quasimetrics, providing an alternative to the classical Temporal-Difference (TD) recursion (sutton2018reinforcement). To accurately represent the optimal value function, wang2023optimal show that the formulated optimization problem must preserve local distances, i.e., the costs of transitions between successive states, while simultaneously pushing apart distant states. Consequently, QRL is framed as the maximization of a quasimetric function where local consistency is enforced through trajectory-based constraints.
@@
Building on these foundations, we propose a novel formulation of QRL that enforces local constraints directly in continuous time rather than through discrete transitions. We argue that for many interesting use cases the discrete-time perspective offers no fundamental benefit beyond practical convenience, as the continuous-time formulation naturally leads to Partial Differential Equations (PDEs). Historically, the difficulties of solving PDEs has limited their role in RL, but recent advances in Physics-Informed Neural Networks (PINNs) (raissi2019physics) have begun to overcome this barrier. By exploiting automatic differentiation, PINNs embed PDE constraints directly into the training objective, enabling neural networks to approximate PDE solutions.
@@
This progress opens the door to a continuous-time perspective on QRL that brings new structural advantages. Most notably, our PDE formulation becomes trajectory-free: rather than relying on short rollouts, it only requires sampling random states and goals from the feasible space. Furthermore, beyond enforcing local consistency, the PDE acts as an implicit regularizer on the value function (giammarino2025physics), improving both learning stability and out-of-distribution estimation accuracy.
@@
    Our main contribution is the introduction of a novel PDE-constrained formulation of QRL that derives continuous local constraints from the Eikonal PDE (noack2017acoustic), leading to Eik-QRL. We provide theoretical guarantees for Eik-QRL and highlight inherent limitations in complex dynamical settings (cf. Section 4).
@@
    Building on Eik-QRL, we propose Eik-Hierarchical QRL (Eik-HiQRL), a hierarchical algorithm that mitigates these limitations while preserving the strengths of Eik-QRL (cf. Section 5). Eik-HiQRL achieves state-of-the-art (SOTA) performance on the OGbench navigation suite (park2024ogbench) and delivers consistent improvements over QRL in manipulation tasks, while also matching the performance of TD-based algorithms.
@@ 2 Related Work
GCRL was introduced to address two longstanding challenges in classical RL (sutton2018reinforcement). The first is the difficulty of designing reward functions, a time consuming process, prone to unintended side effects. Reward shaping (ng1999policy) has been proposed to ease this burden, but it remains subject to reward hacking, where agents exploit loopholes to maximize rewards without achieving the intended behavior (amodei2016concrete; everitt2017reinforcement; everitt2021reward; pan2022effects; di2022goal). Sparse rewards offer an alternative to hand-crafted shaping, yet they introduce complementary challenges due to the weakness of the learning signal. To this end, GCRL provides a natural way to cope with this difficulty by framing tasks in terms of goal reaching, which enables learning directly from sparse success indicators (kaelbling1993learning; schaul2015universal), and has motivated specialized algorithms to improve efficiency under sparse feedback (andrychowicz2017hindsight). A second limitation of classical RL is its focus on single-task settings, which restricts generalization across objectives. By conditioning policies on goals, GCRL extends naturally to multi-goal learning and supports more general-purpose behaviors. Recent advances in this direction include Contrastive RL (eysenbach2022contrastive), state-occupancy matching (ma2022offline), and QRL (wang2023optimal).
@@
Offline RL trains policies from pre-collected datasets without further interaction with the environment (levine2020offline). This paradigm is appealing in domains where online data collection is expensive, risky, or impractical, but it suffers from distributional shift and extrapolation error when policies select actions outside the dataset support. Extending goal conditioning to this setting has given rise to Offline GCRL, which inherits the challenges of both paradigms: agents must generalize to unseen state-goal pairs while learning entirely from fixed data and without corrective feedback. These compounded difficulties have motivated a range of specialized methods (chebotar2021actionable; yang2022rethinking; yang2023essential; mezghani2023learning; sikchi2023smore; park2024hiql; park2024ogbench) aimed at improving stability and generalization. Our work contributes to this line by introducing PDE-based constraints that improve generalization to out-of-distribution state-goal pairs, with particular benefits in large environments and in scenarios that require trajectory stitching.
@@
A key advance in GCRL is recognizing that the optimal goal-conditioned value function corresponds to the shortest feasible path between a state and a goal, and therefore naturally defines a quasimetric (sontag1995abstract; liu2023metric). As mentioned above, QRL (wang2023optimal) builds on this observation by restricting value functions to quasimetric mappings, aligning the hypothesis space with shortest-path goal reaching. This perspective connects to earlier work on quasimetrics in representation learning (sontag1995abstract; pitisinductive; durugkar2021adversarial) and has inspired the development of differentiable architectures (wang2022improved; liu2023metric) that provide universal approximation guarantees for quasimetric parameterized functions. Our work establishes PDE-constrained QRL as a new direction, reformulating value learning in continuous time and revealing structural properties that extend beyond existing formulations.
@@
The continuous-time view of optimal control is traditionally formulated through PDEs, most notably the Hamilton-Jacobi-Bellman (HJB) equation (munos1997convergent; munos1997reinforcement). Although these formulations capture both local and global properties of value functions, they are notoriously difficult to solve in practice, which has long limited their impact in RL. More recently, PINNs (rudy2017data; raissi2019physics; bansal2021deepreach) have leveraged automatic differentiation to embed PDE constraints directly into neural training, enabling progress in areas such as system identification and scientific computing. This paradigm has also been applied to RL: shilova2023revisiting used PINNs to approximate HJB solutions in continuous-time RL, though their results were limited to simple dynamics and low-dimensional state spaces, while other works have employed PDE-inspired regularizers to improve value estimation accuracy in Offline RL (lienenhancing) and Offline GCRL (giammarino2025physics). In contrast, our approach incorporates PDE constraints into the QRL framework, yielding a continuous-time formulation that offers structural benefits not explored in earlier work.
@@
Temporal abstraction is a long-standing idea in RL, where policies are decomposed into high- and low-level components to improve exploration and credit assignment (sutton1999between; vezhnevets2017feudal; nachum2018data; giammarino2021online). Recent work has shown that this structure is particularly effective in GCRL, where subgoal decomposition mitigates the challenges of long horizons (park2024hiql). In our framework, hierarchy is essential, as it both mitigates the practical limitations of Eik-QRL and amplifies its strengths, thereby further improving value learning and overall performance in long-horizon tasks.
@@ 3 Preliminaries
and $\forall x$, $d(x,x)=0$. We denote by $\mathcal{Q}\mathrm{met}(\mathcal{X})$ the space of quasimetrics over $\mathcal{X}\times\mathcal{X}$. QRL is formulated on the observation that, in markovian environments, the optimal goal-conditioned value function, $V^{*}(s,g)=-d^{*}(s,g)$, belongs to the space of quasimetrics. Hence, $V^{*}\in\mathcal{Q}\mathrm{met}(\mathcal{S})$. As a consequence, GCRL can be re-framed as an optimization problem over $\mathcal{Q}\mathrm{met}(\mathcal{S})$. wang2023optimal argue that, in order to capture both local and global relationships between states, this optimization should preserve local distances (i.e., transition costs) between nearby states while maximally spreading out distant ones. This motivates the following formulation:
@@
Figure 1: Diagrams for QRL (wang2020critic). QRL consists of a quasimetric value function parameterized by an Interval Quasimetric Embedding (IQE) model (wang2022improved) and an actor $\pi$.
@@
Here we intentionally reuse the notation $d^{*}(s,g)$, as in the quasimetric setting above, and will later establish a direct connection between the two. Under standard regularity conditions (Assumptions 4.1, 4.2, 4.3 and 4.4), $d^{*}(s(t+\Delta t),g)$ in (2) can be approximated by its first-order Taylor expansion (giammarino2025physics), yielding
@@ 4 Eikonal-Constrained Quasimetric RL
These assumptions are standard optimal-control regularity conditions that ensure a sound HJB formulation (munos1997reinforcement; munos1997convergent; kim2021hamilton). Assumption 4.1 guarantees that the minimization over actions in Eq. (3) is well posed, i.e., the $\arg\min$ in the HJB operator exists. Assumption 4.2 ensures that the system dynamics admit unique and well-defined trajectories since Lipschitz continuity of $f(s,a)$ guarantees existence and uniqueness of ODE solutions (coddington1955theory). This in turn makes the value function and the associated HJB PDE well posed, and provides stability for local expansions and numerical approximations such as $\mathcal{T}(s,a)=s+f(s,a)\Delta t+o(\Delta t)$. Determinism is imposed mainly for analytical convenience and is consistent with prior work (kumar2019stabilizing; wang2023optimal); importantly, our method still applies in stochastic settings, as illustrated by our teleport experiments in Section 6. Finally, Assumption 4.3 ensures finite path costs, and Assumption 4.4 permits point-wise Taylor expansions of value functions.
@@
In the hypothetical case that $a^{*}$ were available, it becomes natural to view the HJB residual in (6) as a surrogate for the discrete local constraint in (4). In practice, if the goal $g$ is chosen as a state visited along the same trajectory as the transition $(s,s^{\prime})$ (andrychowicz2017hindsight), one can reasonably approximate $f(s,a^{*})\approx s^{\prime}-s$, which leads to the following HJB-QRL formulation:
@@
This formulation offers a PINN-style perspective on goal-conditioned value learning, which extends beyond purely additive regularizers explored in prior work (lienenhancing; giammarino2025physics). However, the problem in (7) becomes difficult to optimize in practice, as the inner product $\nabla_{s}d_{\bm{\theta}}(s,g)^{\!\top}(s^{\prime}-s)$ is often ill-conditioned in high-dimensional state spaces, a challenge also noted in many PINN-based approaches to HJB (shilova2023revisiting; bansal2021deepreach). Furthermore, it retains dependence on transition tuples $(s,s^{\prime})$, reflecting QRL’s reliance on trajectories.
As a result, in order to tackle the aforementioned practical challenges, we propose to refine HJB-QRL by imposing explicit assumptions on the system dynamics $f(s,a)$. Specifically, we simplify the HJB PDE to an Eikonal PDE (noack2017acoustic) by leveraging unit-speed, isotropic dynamics $f(s,a)=a$ with $\|a\|\leq 1$. In this setting, it follows that $a^{*}=-\nabla_{s}d(s,g)/\|\nabla_{s}d(s,g)\|$, reducing the local constraint to a unit-slope condition. This yields the following Eikonal-constrained QRL (Eik-QRL) formulation:
@@
A natural concern arises from assuming isotropic dynamics $f(s,a)=a$ with $\|a\|\leq 1$, which may appear restrictive. We emphasize that this choice primarily simplifies the problem numerically, and empirical evidence show that Eik-QRL outperforms HJB-QRL and QRL even under complex dynamics. We acknowledge, however, that this assumption commits us to a class of solutions that might not work optimally for all MDPs and, consequently, envision future work which extend beyond isotropic dynamics, while preserving similar numerical advantages. More broadly, we view both Eik-QRL and HJB-QRL as providing a hybrid regime in which simplified dynamical models impose explicit value-learning constraints that complement model-free RL. Unlike model-based RL (hafner2019learning), where models are mainly used to generate imaginary rollouts, this perspective offers a potential bridge between model-free and model-based approaches.
@@
We now turn to guarantees for Eik-QRL. Formal statements and proofs for these results are deferred to Appendix A. Optimal value recovery guarantees follow from the existence of a quasimetric approximator $d_{\bm{\theta}^{*}}\in\mathcal{Q}\mathrm{met}(\mathcal{S})$ that satisfies $V^{*}(s,g)\;\leq\;d_{\bm{\theta}^{*}}(s,g)\;\leq\;-(1+\epsilon)\,V^{*}(s,g)$ for some small $\epsilon>0$. Given such an approximator, a standard high-probability recovery bound can be derived. However, unlike classical QRL, establishing existence in Eik-QRL is more delicate due to the gradient constraint enforced through the Eik-LRs term in (8). To drive $\epsilon\to 0$ while respecting this constraint, the target $V^{*}(s,g)$ must itself obey a matching gradient regularity (anil2019sorting). The next lemma formalizes this regularity, ensuring that a compatible $d_{\bm{\theta}^{*}}$ exists and closing the argument for optimal value recovery.
@@
Thus, Theorem 4.8 establishes that a universal quasimetric approximator trained with Eik-QRL in (8) recovers $V^{*}(s,g)$ approximately with high probability, under the sufficient conditions of Lemma 4.7. While these conditions provide clean guarantees, they are not strictly required in practice. As shown in our experiments (Section 6), Eik-QRL remains competitive even when $1$-Lipschitz property cannot be verified, as in the antmaze tasks. This observation aligns with the fact that if the target $d^{*}$ is Lipschitz, then approximating it with Lipschitz function, possibly with a different constant, preserves shortest-path geometry up to a positive scaling factor (sethian1996fast), which is typically sufficient for policy gradient methods given their scale-invariance (sutton1999policy; schulman2015trust).
@@
Figure 2: Diagram for Eik-HiQRL. The high-level module consists of a quasimetric model $d^{h}$, parameterized as an IQE (wang2022improved), which is trained to optimize the Eik-QRL objective in Eq. (8). The low-level module comprises a value function $V^{l}$ and a goal representation network $\phi:\mathcal{S}\times\mathcal{G}\to\mathcal{Z}$ (park2024hiql), both instantiated as MLPs and trained via TD-based recursion. Both high-level policy $\pi^{h}$ and the low-level policy $\pi^{l}$ are trained via advantage-weighted regression (peng2019advantage). Full details are provided in Appendix C.

### v2

@@ 1 Introduction
In recent years, Reinforcement Learning (RL) has emerged as a powerful paradigm for solving complex decision-making tasks from experience, achieving success in domains such as game playing (Mnih et al., 2015; Silver et al., 2017), robotics (Tang et al., 2025), autonomous driving (Kiran et al., 2021; Sabouni et al., 2024; Ahmad et al., 2025), and industrial control (Mirhoseini et al., 2020). Despite these achievements, the aforementioned solutions fundamentally rely on human-designed rewards, a process that is notoriously difficult and time consuming. Reward engineering introduces hand-crafted structures, additional hyperparameters, and assumptions about task solutions, which ultimately undermines RL’s promise to transcend human design and ability.
@@
Goal-Conditioned RL (GCRL) provides an elegant way to address this challenge by framing problems in terms of reaching arbitrary goals (Schaul et al., 2015; Andrychowicz et al., 2017). By conditioning policies on goals, GCRL offers a scalable route toward flexible, goal-directed behavior without the need for manually designed reward functions. Beyond this practical advantage, recent work has uncovered a fundamental geometric property of GCRL: the optimal goal-conditioned value function $V^{*}(s,g)$ corresponds to the length of the shortest feasible path from a state $s$ to a goal $g$, and therefore naturally defines a quasimetric (Sontag, 1995; Liu et al., 2023).
@@
This perspective underpins Quasimetric RL (QRL) (Wang et al., 2023), a recently proposed GCRL algorithm that restricts value functions to quasimetric mappings, reducing the hypothesis space from arbitrary functions to a structured subset aligned with goal-reaching tasks. In this framework, GCRL is recast as an optimization problem over quasimetrics, providing an alternative to the classical Temporal-Difference (TD) recursion (Sutton and Barto, 2018). To accurately represent the optimal value function, Wang et al. (2023) show that the formulated optimization problem must preserve local distances, i.e., the costs of transitions between successive states, while simultaneously pushing apart distant states. Consequently, QRL is framed as the maximization of a quasimetric function where local consistency is enforced through trajectory-based constraints.
@@
Building on these foundations, we propose a novel formulation of QRL that enforces local constraints directly in continuous time rather than through discrete transitions. We argue that for many interesting use cases the discrete-time perspective offers no fundamental benefit beyond practical convenience, as the continuous-time formulation naturally leads to Partial Differential Equations (PDEs). Historically, the difficulties of solving PDEs has limited their role in RL, but recent advances in Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) have begun to overcome this barrier. By exploiting automatic differentiation, PINNs embed PDE constraints directly into the training objective, enabling neural networks to approximate PDE solutions.
@@
This progress opens the door to a continuous-time perspective on QRL that brings new structural advantages. Most notably, our PDE formulation becomes trajectory-free: rather than relying on short rollouts, it only requires sampling random states and goals from the feasible space. Furthermore, beyond enforcing local consistency, the PDE acts as an implicit regularizer on the value function (Giammarino et al., 2025), improving both learning stability and out-of-distribution estimation accuracy.
@@
    Our main contribution is the introduction of a novel PDE-constrained formulation of QRL that derives continuous local constraints from the Eikonal PDE (Noack and Clark, 2017), leading to Eik-QRL. We provide theoretical guarantees for Eik-QRL and highlight inherent limitations in complex dynamical settings (cf. Section 4).
@@
    Building on Eik-QRL, we propose Eik-Hierarchical QRL (Eik-HiQRL), a hierarchical algorithm that mitigates these limitations while preserving the strengths of Eik-QRL (cf. Section 5). Eik-HiQRL achieves state-of-the-art (SOTA) performance on the OGbench navigation suite (Park et al., 2024a) and delivers consistent improvements over QRL in manipulation tasks, while also matching the performance of TD-based algorithms.
@@ 2 Related Work
GCRL was introduced to address two longstanding challenges in classical RL (Sutton and Barto, 2018). The first is the difficulty of designing reward functions, a time consuming process, prone to unintended side effects. Reward shaping (Ng et al., 1999) has been proposed to ease this burden, but it remains subject to reward hacking, where agents exploit loopholes to maximize rewards without achieving the intended behavior (Amodei et al., 2016; Everitt et al., 2017; 2021; Pan et al., 2022; Di Langosco et al., 2022). Sparse rewards offer an alternative to hand-crafted shaping, yet they introduce complementary challenges due to the weakness of the learning signal. To this end, GCRL provides a natural way to cope with this difficulty by framing tasks in terms of goal reaching, which enables learning directly from sparse success indicators (Kaelbling, 1993; Schaul et al., 2015), and has motivated specialized algorithms to improve efficiency under sparse feedback (Andrychowicz et al., 2017). A second limitation of classical RL is its focus on single-task settings, which restricts generalization across objectives. By conditioning policies on goals, GCRL extends naturally to multi-goal learning and supports more general-purpose behaviors. Recent advances in this direction include Contrastive RL (Eysenbach et al., 2022), state-occupancy matching (Ma et al., 2022), and QRL (Wang et al., 2023).
@@
Offline RL trains policies from pre-collected datasets without further interaction with the environment (Levine et al., 2020). This paradigm is appealing in domains where online data collection is expensive, risky, or impractical, but it suffers from distributional shift and extrapolation error when policies select actions outside the dataset support. Extending goal conditioning to this setting has given rise to Offline GCRL, which inherits the challenges of both paradigms: agents must generalize to unseen state-goal pairs while learning entirely from fixed data and without corrective feedback. These compounded difficulties have motivated a range of specialized methods (Chebotar et al., 2021; Yang et al., 2022; 2023; Mezghani et al., 2023; Sikchi et al., 2023; Park et al., 2024b; a) aimed at improving stability and generalization. Our work contributes to this line by introducing PDE-based constraints that improve generalization to out-of-distribution state-goal pairs, with particular benefits in large environments and in scenarios that require trajectory stitching.
@@
A key advance in GCRL is recognizing that the optimal goal-conditioned value function corresponds to the shortest feasible path between a state and a goal, and therefore naturally defines a quasimetric (Sontag, 1995; Liu et al., 2023). As mentioned above, QRL (Wang et al., 2023) builds on this observation by restricting value functions to quasimetric mappings, aligning the hypothesis space with shortest-path goal reaching. This perspective connects to earlier work on quasimetrics in representation learning (Sontag, 1995; Pitis et al., ; Durugkar et al., 2021) and has inspired the development of differentiable architectures (Wang and Isola, ; Liu et al., 2023) that provide universal approximation guarantees for quasimetric parameterized functions. Our work establishes PDE-constrained QRL as a new direction, reformulating value learning in continuous time and revealing structural properties that extend beyond existing formulations.
@@
The continuous-time view of optimal control is traditionally formulated through PDEs, most notably the Hamilton-Jacobi-Bellman (HJB) equation (Munos, 1997; Munos and Bourgine, 1997). Although these formulations capture both local and global properties of value functions, they are notoriously difficult to solve in practice, which has long limited their impact in RL. More recently, PINNs (Rudy et al., 2017; Raissi et al., 2019; Bansal and Tomlin, 2021) have leveraged automatic differentiation to embed PDE constraints directly into neural training, enabling progress in areas such as system identification and scientific computing. This paradigm has also been applied to RL: Shilova et al. (2023) used PINNs to approximate HJB solutions in continuous-time RL, though their results were limited to simple dynamics and low-dimensional state spaces, while other works have employed PDE-inspired regularizers to improve value estimation accuracy in Offline RL (Lien et al., 2024) and Offline GCRL (Giammarino et al., 2025). In contrast, our approach incorporates PDE constraints into the QRL framework, yielding a continuous-time formulation that offers structural benefits not explored in earlier work.
@@
Temporal abstraction is a long-standing idea in RL, where policies are decomposed into high- and low-level components to improve exploration and credit assignment (Sutton et al., 1999b; Vezhnevets et al., 2017; Nachum et al., 2018; Giammarino and Paschalidis, 2021). Recent work has shown that this structure is particularly effective in GCRL, where subgoal decomposition mitigates the challenges of long horizons (Park et al., 2024b). In our framework, hierarchy is essential, as it both mitigates the practical limitations of Eik-QRL and amplifies its strengths, thereby further improving value learning and overall performance in long-horizon tasks.
@@ 3 Preliminaries
and $\forall x$, $d(x,x)=0$. We denote by $\mathcal{Q}\mathrm{met}(\mathcal{X})$ the space of quasimetrics over $\mathcal{X}\times\mathcal{X}$. QRL is formulated on the observation that, in markovian environments, the optimal goal-conditioned value function, $V^{*}(s,g)=-d^{*}(s,g)$, belongs to the space of quasimetrics. Hence, $V^{*}\in\mathcal{Q}\mathrm{met}(\mathcal{S})$. As a consequence, GCRL can be re-framed as an optimization problem over $\mathcal{Q}\mathrm{met}(\mathcal{S})$. Wang et al. (2023) argue that, in order to capture both local and global relationships between states, this optimization should preserve local distances (i.e., transition costs) between nearby states while maximally spreading out distant ones. This motivates the following formulation:
@@
Figure 1: Diagrams for QRL (Wang et al., 2020). QRL consists of a quasimetric value function parameterized by an Interval Quasimetric Embedding (IQE) model (Wang and Isola, ) and an actor $\pi$.
@@
Here we intentionally reuse the notation $d^{*}(s,g)$, as in the quasimetric setting above, and will later establish a direct connection between the two. Under standard regularity conditions (Assumptions 4.1, 4.2, 4.3 and 4.4), $d^{*}(s(t+\Delta t),g)$ in (2) can be approximated by its first-order Taylor expansion (Giammarino et al., 2025), yielding
@@ 4 Eikonal-Constrained Quasimetric RL
These assumptions are standard optimal-control regularity conditions that ensure a sound HJB formulation (Munos and Bourgine, 1997; Munos, 1997; Kim et al., 2021). Assumption 4.1 guarantees that the minimization over actions in Eq. (3) is well posed, i.e., the $\arg\min$ in the HJB operator exists. Assumption 4.2 ensures that the system dynamics admit unique and well-defined trajectories since Lipschitz continuity of $f(s,a)$ guarantees existence and uniqueness of ODE solutions (Coddington and Levinson, 1955). This in turn makes the value function and the associated HJB PDE well posed, and provides stability for local expansions and numerical approximations such as $\mathcal{T}(s,a)=s+f(s,a)\Delta t+o(\Delta t)$. Determinism is imposed mainly for analytical convenience and is consistent with prior work (Kumar et al., 2019; Wang et al., 2023); importantly, our method still applies in stochastic settings, as illustrated by our teleport experiments in Section 6. Finally, Assumption 4.3 ensures finite path costs, and Assumption 4.4 permits point-wise Taylor expansions of value functions.
@@
In the hypothetical case that $a^{*}$ were available, it becomes natural to view the HJB residual in (6) as a surrogate for the discrete local constraint in (4). In practice, if the goal $g$ is chosen as a state visited along the same trajectory as the transition $(s,s^{\prime})$ (Andrychowicz et al., 2017), one can reasonably approximate $f(s,a^{*})\approx s^{\prime}-s$, which leads to the following HJB-QRL formulation:
@@
This formulation offers a PINN-style perspective on goal-conditioned value learning, which extends beyond purely additive regularizers explored in prior work (Lien et al., 2024; Giammarino et al., 2025). However, the problem in (7) becomes difficult to optimize in practice, as the inner product $\nabla_{s}d_{\bm{\theta}}(s,g)^{\!\top}(s^{\prime}-s)$ is often ill-conditioned in high-dimensional state spaces, a challenge also noted in many PINN-based approaches to HJB (Shilova et al., 2023; Bansal and Tomlin, 2021). Furthermore, it retains dependence on transition tuples $(s,s^{\prime})$, reflecting QRL’s reliance on trajectories.
As a result, in order to tackle the aforementioned practical challenges, we propose to refine HJB-QRL by imposing explicit assumptions on the system dynamics $f(s,a)$. Specifically, we simplify the HJB PDE to an Eikonal PDE (Noack and Clark, 2017) by leveraging unit-speed, isotropic dynamics $f(s,a)=a$ with $\|a\|\leq 1$. In this setting, it follows that $a^{*}=-\nabla_{s}d(s,g)/\|\nabla_{s}d(s,g)\|$, reducing the local constraint to a unit-slope condition. This yields the following Eikonal-constrained QRL (Eik-QRL) formulation:
@@
A natural concern arises from assuming isotropic dynamics $f(s,a)=a$ with $\|a\|\leq 1$, which may appear restrictive. We emphasize that this choice primarily simplifies the problem numerically, and empirical evidence show that Eik-QRL outperforms HJB-QRL and QRL even under complex dynamics. We acknowledge, however, that this assumption commits us to a class of solutions that might not work optimally for all MDPs and, consequently, envision future work which extend beyond isotropic dynamics, while preserving similar numerical advantages. More broadly, we view both Eik-QRL and HJB-QRL as providing a hybrid regime in which simplified dynamical models impose explicit value-learning constraints that complement model-free RL. Unlike model-based RL (Hafner et al., 2019), where models are mainly used to generate imaginary rollouts, this perspective offers a potential bridge between model-free and model-based approaches.
@@
We now turn to guarantees for Eik-QRL. Formal statements and proofs for these results are deferred to Appendix A. Optimal value recovery guarantees follow from the existence of a quasimetric approximator $d_{\bm{\theta}^{*}}\in\mathcal{Q}\mathrm{met}(\mathcal{S})$ that satisfies $V^{*}(s,g)\;\leq\;d_{\bm{\theta}^{*}}(s,g)\;\leq\;-(1+\epsilon)\,V^{*}(s,g)$ for some small $\epsilon>0$. Given such an approximator, a standard high-probability recovery bound can be derived. However, unlike classical QRL, establishing existence in Eik-QRL is more delicate due to the gradient constraint enforced through the Eik-LRs term in (8). To drive $\epsilon\to 0$ while respecting this constraint, the target $V^{*}(s,g)$ must itself obey a matching gradient regularity (Anil et al., 2019). The next lemma formalizes this regularity, ensuring that a compatible $d_{\bm{\theta}^{*}}$ exists and closing the argument for optimal value recovery.
@@
Thus, Theorem 4.8 establishes that a universal quasimetric approximator trained with Eik-QRL in (8) recovers $V^{*}(s,g)$ approximately with high probability, under the sufficient conditions of Lemma 4.7. While these conditions provide clean guarantees, they are not strictly required in practice. As shown in our experiments (Section 6), Eik-QRL remains competitive even when $1$-Lipschitz property cannot be verified, as in the antmaze tasks. This observation aligns with the fact that if the target $d^{*}$ is Lipschitz, then approximating it with Lipschitz function, possibly with a different constant, preserves shortest-path geometry up to a positive scaling factor (Sethian, 1996), which is typically sufficient for policy gradient methods given their scale-invariance (Sutton et al., 1999a; Schulman et al., 2015).
@@
Figure 2: Diagram for Eik-HiQRL. The high-level module consists of a quasimetric model $d^{h}$, parameterized as an IQE (Wang and Isola, ), which is trained to optimize the Eik-QRL objective in Eq. (8). The low-level module comprises a value function $V^{l}$ and a goal representation network $\phi:\mathcal{S}\times\mathcal{G}\to\mathcal{Z}$ (Park et al., 2024b), both instantiated as MLPs and trained via TD-based recursion. Both high-level policy $\pi^{h}$ and the low-level policy $\pi^{l}$ are trained via advantage-weighted regression (Peng et al., 2019). Full details are provided in Appendix C.